what is the slope of a line ( -3 ,2) (3 -2)
Answer by Cromlix(4381)

Equation for slope of a line
= y2 - y1/x2 - x1
= -2 -2/3-(-3)
= -4/6
= -2/3
Hope this helps.
:-)